OpenStudy (anonymous):

R(x)=7x+14/5-x please help! looking for asymptote

OpenStudy (anonymous):

I found x=5 but not sure

OpenStudy (aonz):

5 is vertical asymptote

OpenStudy (anonymous):

On the same equation: The horizontal asymptote is defined by the function: x=7/5 x = -7 Rx=7/5 R(x) = -7

OpenStudy (anonymous):

I think you are correct. Think about what happens as x gets closer and closer to 5. The numerator gets divided by a smaller and smaller number the closer x gets to 5. What happens when you take some constant and divide it by something really, /really/ small?

OpenStudy (anonymous):

I see...yes it is the vertical (x=5) they asking for horizontal...

OpenStudy (aonz):

sub those values in y and find which one will not be possible

OpenStudy (anonymous):

oh, for the horizontal I think you might have some rules in your book to guide you. Some like if n >m, or n = m, or n < m, with n and me being the degree of the polynomials in the numerator and denominator.

OpenStudy (aonz):

if you do you would find that R(x) = -7 is impossible

OpenStudy (anonymous):

since the degree (of one) is the same in the top and bottom of your fraction, you have a case of n = m. If I remember, that is just the ration of the coeffiecients, 7/(-1).

OpenStudy (anonymous):

I would have answered that h.a. is y = -7 also. So if that is incorrect I guess I would have got it wrong.

OpenStudy (anonymous):

Ok thank you very much! so should i answer This has no solution, so the denominator is never zero. Then the domain is "all x".?

OpenStudy (aonz):

so you want to find the domain of this?

OpenStudy (anonymous):

yes,correct

OpenStudy (aonz):

well, domain is x all real except 5

OpenStudy (anonymous):

if x = 5 is a v.a. then you have to exclude it from your domain. Unless you figured out how to divide by zero.

OpenStudy (aonz):

as x cannot be 5 since denominator cannot be 0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think you've got everything you need: h.a. is y = -7, v.a. is x = 5. Domain is {x| x is not equal to 5} or maybe like this, (-infinity, 5) U (5, infinity).

OpenStudy (aonz):

i dont think she wants to sketch this lol

OpenStudy (anonymous):

lol. personally, I would have just put it in my graphing calculator unless I needed to show the work.

OpenStudy (anonymous):

i did sketch it!! :)

OpenStudy (anonymous):

i do have to show my work, and my graph is a T85...old and i don't get how to use it. but thank you for the help

OpenStudy (aonz):

check if you got your sketch right :P

OpenStudy (anonymous):

this is a great link, thanks!

OpenStudy (aonz):

btw that website can solve any algebra question or calculus question with working out... Sadly if you want working out its not free

OpenStudy (anonymous):

well it does help..